silver has two naturally occurring isotopes, one of mass 106.91 amu and the other of mass 108.90amu. FInd the fractional abundances for these two isotopes. The atomic weight is 107.87 amu
1
Expert's answer
2017-06-10T05:56:12-0400
Notice that the abundance of Ag-107 (106.91 amu) is assigned “x” and the Ag-109 (108.90 amu) is “1 – x”.
106.91x + 108.90(1 – x) = 107.87
106.91x + 108.90 – 108.90x = 107.87
1.99x = 1.03
x = 0.518
The fractional abundance of second isotope is 1 – 0.518 = 0.482.

Answer:
The fractional abundance of Ag-107 is 0.518,
the fractional abundance of Ag-109 is 0.482.